OpenStudy (anonymous):

The parabola y=ax^2+bx+3 has a vertex (-2,-1) find a and b

OpenStudy (anonymous):

Some one help!

OpenStudy (akashdeepdeb):

The standard equation of parabola = y = ax 2 + bx + c And the vertex, x co-ordinate = x = -b/2a okay? And we have a point on the parabola = (x,y) = (-2,-1) And so -1 = 4a - 2b + 3 or 4a - 2b = -4 ---- 1 x = -b/2a -2(2a) = -b -4a = -b --- 2 Use the 2 equations understood? :)

OpenStudy (akashdeepdeb):

Refer to this if you have any doubt! http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php

OpenStudy (akashdeepdeb):

Are you getting me?

OpenStudy (anonymous):

ok thanks!

OpenStudy (akashdeepdeb):

:)